a body cools from 67 to 37 in t secs at surrounding temp of 27. if the surrounding temp is 7, how much time would it take?

Grade:12th pass

2 Answers

Khimraj
5 years ago
ACco
Khimraj
5 years ago
According to newtons law of cooling
log(T1 -T)/(T2 – T) = Kt
log(67 – 27)/(37-27) = Kt
log4 = Kt
t = 2log2/K
and final condition is
log(67 – 7)/(37 – 7) = Kt’
log2 = Kt’
t’ = log2/K = t/2.
Hope it clears.